# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v5

@@ Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. Therefore, randomly initialized models are often unnecessarily large, even when approximating simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to exhibit a number of linear regions exponential in depth. Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts. We further demonstrate how to extend our approach to multidimensional and non-convex functions, with similar benefits observed.
@@ 1 Introduction
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin & Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although preliminary numerical demonstrations in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. In one dimension, where we initially focus, our work is a generalization of the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions, and then converts that result back into a matrix representation. The goal of this method is not to re-express the full set of neural network weights, but to constrain them into efficient patterns.
@@ 2.1 Function Approximation
The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)). Our reparameterization strategy focuses on non-symmetric triangle waves. The location on $(0,1)$ of a triangular function’s peak acts as an alternative parameterization of the function, instead of its weight matrix representation. Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions. Furthermore, this parameterization is trainable. By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions.
The dilated triangular waveforms produced in this manner are not particularly useful on their own. Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit). But these problems can be rectified by taking a sum over the layers of a network. Yarotsky (2017) and Liang & Srikant (2016) construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$. Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
@@ 2.2 Neural Network Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. As alluded to in Section 2.1, our initialization algorithm is simply to associate a nonsymmetric triangle function with each layer, and then to randomize the location of each triangle’s peak in $(0,1)$; any choice produces $2^{d}$ regions. This yields a reparameterization of the ReLU network in terms of the triangle peak locations, $a_{i}$, rather than in terms of the network’s weights. The pretraining step of our algorithm then involves learning and updating these parameters. This is done by setting individual neuron weights in accordance with the procedures described in this section, and then backpropagating the gradients to update the triangle peak parameters. Pseudocode for our entire algorithm is provided in Appendix A.1.
@@
where $10<a_{i}<$ 1. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$. Each layer considered in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise linear functions:
@@
Figure 2: On the right is a network representation of a triangle function. The middle shows that triangle function as a hidden layer of a network. The one-dimensional input and output of a triangle function is converted into shared weights. A full pretraining network is assembled on the left.
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
There are two other neurons in the pretraining network’s layers, each with a specific role. The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer. This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions). If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer. We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially in order to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
@@ 3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters. This would be poorly predictive of unseen data points generated from a smooth curve. Our main mathematical result addresses this issue by forcing differentiability of the network output. We do so by making the sum of the derivatives of the triangle waves converge. For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them. For a broader mathematical discussion, as well as sufficient conditions for differentiability in the limit, see the appendix.
@@ 4 One-Dimensional Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark and demonstrate results on more difficult functions in Section 5, Section 6, and Appendix A.10. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3). We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is to see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix for a brief analysis of different learning rates). Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations using the construction in Section 3, so that the effects of each training regimen are directly comparable. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we focus for the moment on one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one-dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.

### v6

@@ Abstract
In a neural network with ReLU activations, the number of piecewise linear regions in the output can grow exponentially with depth. However, this is highly unlikely to happen when the initial parameters are sampled randomly, which therefore often leads to the use of networks that are unnecessarily large. To address this problem, we introduce a novel parameterization of the network that restricts its weights so that a depth $d$ network produces exactly $2^{d}$ linear regions at initialization and maintains those regions throughout training under the parameterization. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts. We further demonstrate how to extend our approach to multidimensional and non-convex functions, allowing it to replace the dense layers in other networks; preliminary improvements are shown for image classification on CIFAR-10 and ImageNet.
Machine Learning, ICML
@@ 1 Introduction
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Indeed, Hanin & Rolnick (2019) show a rather disappointing bound for randomly initialized networks: the average number of linear regions formed at initialization is invariant with depth. Whether using a deep or shallow architecture, only the total number of neurons matters. They further observed that gradient descent has a difficult time creating new activation regions so that their bounds approximately held after training. As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see 5). Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
The present work presents mathematical algorithms that avoid the limitations of random initialization. Our contributions include:
  • A reparameterization of a 4-neuron-wide, depth $d$ ReLU network that constrains it to initialize with $2^{d}$ activation regions over the input domain (Section 3)
  • A novel pretraining strategy, which enforces the existence of $2^{d}$ activation regions during optimization to produce its initial solution (Sections 3, 3.1)
  • Numerical results for one-dimensional test cases that yield orders-of-magnitude improvements in network performance (Section 4)
  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Section 5; Appendices A.10A.11)
While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and quickly outpacing such networks’ performance on more practical test cases such as ImageNet (Appendix A.11). We emphasize that while the preliminary mathematical analysis in this work pertains to a specific ReLU construction, this construction can be repeated and combined in useful ways to obtain performance improvements on larger networks (Section 5). We conclude in Section 6 with discussions on extending the mathematical framework of our method to arbitrary ReLU networks and other types of networks, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology for ReLU networks, but its development stems from a more abstract idea of an exponentially accurate approximation to the function $x^{2}$, which can be encoded into a ReLU network. Our work generalizes this construction into a trainable family of differentiable convex one-dimensional functions (see Appendix A.3).
@@ 2.1 Function Approximation
The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)). Our reparameterization strategy (Section 3) focuses on non-symmetric triangle waves.
The dilated triangular waveforms produced in this manner are not particularly useful on their own. Their oscillations quickly become excessively rapid, and their derivatives do not exist everywhere (especially in the infinite-depth limit). But these problems can be rectified by taking a sum over the layers of a network. Yarotsky (2017) and Liang & Srikant (2016) construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$. Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
@@ 2.2 Neural Network Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are randomly sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate, the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
We begin this section by describing our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth. We focus on architecting the weights of a 4-neuron-wide, depth $d$ ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
As alluded to in Section 2.1, the overall strategy of our reparameterization and initialization algorithm is to associate a nonsymmetric triangle function with each layer of the network. This yields a reparameterization of the ReLU network in terms of the locations of each triangle’s peak on $(0,1)$, $a_{i}$, rather than in terms of the network’s weights. Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions. Furthermore, this parameterization is trainable. By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions. Pseudocode for our entire algorithm is provided in Appendix A.1.
@@
where $10<a_{i}<$ 1. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$. Each layer considered in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise constant functions:
@@
Figure 2: On the top left is a network representation of a triangle function. The top right shows that triangle function as a hidden layer of a network. The one-dimensional input and output of a triangle function is converted into shared weights. A full pretraining network is assembled on the bottom.
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$ shape, but this is unnecessarily deep and irregular. We can replicate the one-dimensional function composition in the hidden layers of our network (see Figure 2 top right). Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons possess identical slopes in regions where both neurons are active.
There are two other neurons in the pretraining network’s layers, each with a specific role. The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer. This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions). If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer. We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
3.1 Pretraining and Overall Algorithm
We imagine the process of optimizing neural network parameters as three phases: (1) reparameterization and initialization, where a network’s weights are initially set; (2) pretraining, where the network is trained under that reparameterization; and (3) training, where the network is trained using the network’s weights directly. Our pretraining algorithm acts as a preconditioner for or guide to the loss landscape, and when an optimizer is close to a minimum, the exponential expressivity of the network may no longer need to be guaranteed to ensure efficient neuron usage.
Our pretraining algorithm is thus to simply use an optimizer like gradient descent to train the network under the reparameterization described above. Explicit formulas for updating the reparameterized weights, and pseudocode for our whole algorithm, are listed in Appendix A.1.
@@ 3.2 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters. This would be poorly predictive of unseen data points generated from a smooth curve. Our main mathematical result addresses this issue by forcing differentiability of the network output. We do so by making the sum of the derivatives of the triangle waves converge. For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them. Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3.
@@ 4 One-Dimensional Experiments
Since our network construction applies directly to one dimensional convex functions, this will be the focus of this section (we cover more difficult functions in Sections 56 and A.10A.11) The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. We also train several permutations of our reparameterized networks, which validate that our pretraining and our differentibility constraints indeed facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix A.9 for a brief analysis of different learning rates). Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks using triangle based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). In our tables, “pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization. “differentiability not enforced” is a network that initializes with our parameterization, but then trains the scales of its triangle waves as an extra degree of freedom (so that the final output is not necessarily differentiable). “differentiability enforced” corresponds to the main variant of our method as we have presented it thus far. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. This set of experiments evaluates the predictive capacities of the networks on unseen data.